When 732.5 J of heat are added to 102G of a solid substance the temperature of the solid increases by 48.8 K what is the specific heat capacity of the solid


1
Expert's answer
2021-03-08T06:11:41-0500

"Q = mcΔT \\

c = \frac{Q}{mΔT} \\

= \frac{732.5}{102 \times 48.8} \\

= 0.147\;J/gK"

Answer: 0.147 J/gK